A rectangle is twice as long as it is wide. If both it's dimensions are increased by 4 m it's area is increased by 88 m^2. Find the dimensions of the original rectangle.
Answer by oberobic(2304)

L = 2W
.
(L+4)(W+4) = L*W + 88
.
Substitute L = 2W
.
(2W +4)(W +4) = 2W*W +88
.
2W^2 +8W + 4W + 16 = 2W^2 + 88
.
2W^2 +8W + 4W -2W^2 = 88 -16
.
12W = 72
.
W = 6 m = original width
.
L = 2W = 12 m = original length
.
Check the solution to make sure it's the answer.
.
L*W = 72 = area of the original rectangle
.
(L+4)(W+4) = 16*10 = 160 m^2
.
160 -72 = 88 m^2
.
Correct.
.
Done.